OpenStudy (anonymous):

You drop a rock off a bridge. The rock’s height, h (in feet above the water), after t seconds is modeled by h = -16t2 + 541. What is the height of the rock after 2 seconds?

OpenStudy (nurali):

h = -16t^2 + 541 h(2) = - 16(4) + 541 = -64 + 541 = 477 The height of the rock is 477 ft.